OpenStudy (weaver):

What is another way to solve this equation? @will.h 2n + 3 = 24 Given n + 3 = 12 Division Property of Equality n = 11 Subtraction Property of Equality

OpenStudy (countryboyswag):

What do you think? @Weaver

OpenStudy (weaver):

idk @countryboyswag im not good at reasoning and proof

OpenStudy (countryboyswag):

Well, Does one seem to make more sense with one of the answers they have?

OpenStudy (weaver):

there is no answer choices

OpenStudy (will.h):

2n + 3 = 24 At first we can isolate 2b by subtracting 3 from both sides 2n = 24 -3 2n = 21 Now divide both sides by 2 n= 21/2 n= 10.5 I think there's error in the 1st solution

OpenStudy (countryboyswag):

I agree with Will, must be an error.

OpenStudy (will.h):

Am quite certain that the original equation indicates an error If you get to substitute 11 instead of x in the original equations the sides must be equal to each other if the solution was real so 2n + 3 = 24 if we say x= 11 2 * 11 + 3 = 24 25= 24 Not equal. however in my solution they are 2n + 3 = 24 when x= 10.5 2 * 10.5 + 3 = 24 24 = 24